OpenStudy (tal_tal):

Multiply (−4) × (−2) × (−5). −40 −8 8 40

OpenStudy (anonymous):

A

OpenStudy (tal_tal):

Thank you! Is that that right tho?

OpenStudy (anonymous):

Use a calculator? X) If not, 4 x 2=8 8x5=40. There are 3 negatives being multiplied which equals a negative So, -40 :)

OpenStudy (tal_tal):

Okay thank you guys can you help me with one more

OpenStudy (anonymous):

Yes... -4x-2 is 8, 8x-5 is -40

OpenStudy (anonymous):

sure

OpenStudy (tal_tal):

Paul purchased 3 video games for $17.25 each. Which expression is the best choice to help him determine the total amount of money he spent? (3) × (−17.25) = −51.75; He spent $51.75. (−3) × (17.25) = −51.75; He spent $51.75. (3) × (17.25) = 51.75; He spent $51.75. (−3) × (−17.25) = 51.75; He spent $51.75.

OpenStudy (anonymous):

C

OpenStudy (tal_tal):

Thank you very much!!!